OpenStudy (happykiddo):

The volume of a hexagonal prism is 6square root 3 ft3. What is the height of the prism if the base area is 2square root 3 ft3? Answer 3square root 3ft 12 ft 18square root 3 ft 3 ft

OpenStudy (campbell_st):

for a prism V = base area x height so \[6\sqrt{3} = 2\sqrt{3} \times\] \[h = 6\sqrt{3}/2\sqrt{3}\] just evaluate to find the height

OpenStudy (campbell_st):

1st equations should be \[6\sqrt{3} = 2\sqrt{3} \times h\]

OpenStudy (happykiddo):

i got 3.00000000000001

OpenStudy (happykiddo):

so last answer right

OpenStudy (campbell_st):

yes the answer is 3 feet